It is reported that 15% of people who come into contact with a person infected with strep throat contract the disease. How might we use the random-number table to estimate the probability that at least one child in a three-child family will catch the disease, given that each child has come into contact with an infected person?
Answer by ikleyn(52776)

.

"At least one" is the complement event to "no one".


It gives the ANSWER


    P(at least one) = 1 - P(no one) = 1 - (1-0.15)^3 = 1 - 0.85^3 = 0.385875 (rounded).

Solved, with short and complete explanation.